A group of private investors purchased a condominium complex for $3 million. They made an initial down payment of 15% and obtained financing for the balance. If the loan is to be amortized over 11 years at an interest rate of 9.1%/year compounded quarterly, find the required quarterly payment. (Round your answer to the nearest cent.)
$
Answer #1

This question requires application of PV of annuity, according to which

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

PV = Loan Amount = $3,000,000 * (1 - 15%) = $2,550,000

r = 9.1%/4 = 2.275% (quarterly)

n = 11 * 4 = 44 quarters

2,550,000 = P* ,1-(1 + 0.02275) -44 0.02275

2,550,000 = P * 27.619465

P = $92,326.2